OpenStudy (anonymous):

What are the solutions of 2x^2 = 24x?

OpenStudy (anonymous):

x = 0, x = 12 x = 0, x = −12 x = 2, x = 12 x = 2, x = −12

OpenStudy (ajprincess):

2x^2=24x Divide by 2 both sides. then u will get x^2=12x x^2-12x=0 Try to factorise this. Pls tell me vat u get?

OpenStudy (anonymous):

x = 0, x = 12 that wat i got @ajprincess

OpenStudy (ajprincess):

ya u r right:)

OpenStudy (anonymous):

thnx

OpenStudy (ajprincess):

welcome:)